A body covers 10 m at 4 m/s and covers rest of the journey at uniform speed. If the average speed of the body to cover the complete journey of 15 m is 5m/s2, calculate its speed after covering 10 m.



1
Expert's answer
2021-02-23T10:05:34-0500

I assume that there is a mistake in units and it should be "a = 4 m\/s^2" and "v_{av} = 5 m\/s" . Then, by definition of average speed, the time required to complete the journey is:


"t_{total} = \dfrac{15m}{5m/s} = 3s"

Let's denote the speed after covering "d_1 = 10m" as "v_2". Then, by definition of the speed of motion with uniform acceleration, obtain:


"v_2 = v_1 + at_1"

where "v_1" is the speed at the very beginning of the motion, and "t_1" is the time required to cover "d_1".

The distance covered under the uniform acceleration is given as follows:


"d_1 = v_1t_1 + \dfrac{at_1^2}{2}"

Substituting here "v_1" from the previous equation, obtain:


"d_1 = (v_2 - at_1)t_1 + \dfrac{at_1^2}{2} = v_2t_1 - \dfrac{at_1^2}{2}"

Since the second part of the journey "d_2 = 5m" was covered with uniform spee, then have:


"v_2 = \dfrac{d_2}{t_2} = \dfrac{d_2}{t_{total} - t_1}"

where "t_2 = t_{total} - t_1" is the time required to cover "d_2" with uniform speed "v_2". Expressing "t_1" from the last equation and substituting it into the previous one, obtain:


"t_1 = t_{total} - \dfrac{d_2}{v_2}\\
d_1 = v_2\left( t_{total} - \dfrac{d_2}{v_2} \right) - \dfrac{a}{2}\left( t_{total} - \dfrac{d_2}{v_2} \right)^2=\\
=v_2t_{total} - d_2 - \dfrac{a}{2}\left( t_{total}^2 - 2\dfrac{t_{total}d_2}{v_2} + \dfrac{d_2^2}{v_2^2} \right)"

Multiplying by "v_2^2", obtain:


"v_2^3t_{total} - v_2^2\left( d_1 + d_2 + \dfrac{at_{total}^2}{2} \right) - v_2at_{total}d_2 + \dfrac{ad_2^2}{2} = 0\\
v_2^3\cdot 3 - v_2^2\left( 10 +5 + \dfrac{4\cdot 3^2}{2} \right) - v_2\cdot 4\cdot 3\cdot 5+ \dfrac{4\cdot 5^2}{2} = 0\\
3v_2^3 -33 v_2^2 - 60v_2+ 50 = 0"

Solving this equation, obtain:


"v_2 \approx 12.49\space \text{or} \space v_2 \approx 0.63\space \text{or} \space v_2 \approx-2.12"

The speed was assumed to be positive, thus, we can ignore the last value. The second value does not fit as well, since then it is required to spend "t_2 = d_2/v_2 \approx 8s" to cover the second part, which is more than total time of journey "t_{total}". Thus, the only value remain is


"v_2 \approx 12.49m/s"

Answer. 12.49 m/s.